OpenStudy (loser66):

Prove \(lim_{n-->\infty} \dfrac{1}{n+1} =0\) Please help

OpenStudy (loser66):

We know that \[lim_{n\rightarrow \infty}\dfrac{1}{n}=0\] but how to link them together?

OpenStudy (anonymous):

Make a comparison: \[\frac{1}{n+1}\le\frac{1}{n}\] If you can prove the limit for \(\dfrac{1}{n}\), you can prove the the other limit.

OpenStudy (loser66):

I am allowed to use the fact that lim (1/n) =0

OpenStudy (loser66):

Can I use squeeze theorem? but if using it, I have to prove that lim (1/n+1) exists first.

OpenStudy (anonymous):

What kind of definition are you using? The \(\epsilon-N\) for sequences or \(\epsilon-\delta\) for functions?

OpenStudy (loser66):

I am sorry, I don't get you.

OpenStudy (loser66):

You are talking about epsilon? I have to use limit theorem to prove it

OpenStudy (loser66):

I can't ( not allowed) use definition of limit.

OpenStudy (anonymous):

The \(\epsilon-N\) definition of the limit applies to sequences. It says \[\lim_{n\to\infty}a_n=L\] means there exists \(N\) such that \(N\ge n\) implies \(|a_n-L|<\epsilon\) for any \(\epsilon>0\). In this case, you would want to find such an \(N\) that makes \(\left|\dfrac{1}{n+1}\right|<\epsilon\). You can show this easily using the fact that \(n+1>n\) for all \(n\), which gives \[\frac{1}{n+1}<\frac{1}{n}<\epsilon~~\implies~~N=\left\lceil\frac{1}{\epsilon}\right\rceil\] where \(\lceil x\rceil\) denotes the ceiling function, which gives the smallest integer greater than \(x\).

OpenStudy (loser66):

I got you. Thank you :)

OpenStudy (anonymous):

You're welcome!

OpenStudy (freckles):

You can also do this, right? Proof: \[\text{ Let } \epsilon>0 \text{. Set } \\ M=max\left\{ -1, \frac{1-\epsilon}{\epsilon} \right\} \\ \text{ If } n>M \text{, then } n>-1 \text{ and } n> \frac{1- \epsilon}{\epsilon} \text{, } \\ \\ \text{ so } n > \frac{1- \epsilon}{\epsilon} \\ \epsilon n > 1- \epsilon \\ \epsilon n +\epsilon > 1 \\ \epsilon(n+1)>1 \\ \epsilon>\frac{1}{n+1 } \\ |\frac{1}{n+1}-0| < \epsilon \]